Find the length of X in the triangle. Express in simplified radical form:
Here is a drawing of the triangle:
=
= =
= =
12 = = X
= =
= =
= =
= ====
= = =
= = =
=======================
8
I do not know the process of figuring this out, can anyone help me understand the process and how to get the answer please, thank you.

Answer by Nate(3500)

You have a right triangle with the legs: 8 and x
You have a right triangle with the hypotenuse: 12
a^2 + b^2 = c^2
8^2 + x^2 = 12^2
64 + x^2 = 144
x^2 = 80
x = 4*sqrt(5)